OpenStudy (precalchelp):

f(x)=x^(1+cubedrootof4) find the derivative

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ f(x)=x^{1+\sqrt[3]{4}} }\)

OpenStudy (solomonzelman):

In this case there is no reason for you not to apply the power rule. Not kidding, just the power rule :)

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{d}{dx}~x^n=nx^{n-1} }\) \(\Large\color{blue}{ n=1+\sqrt[3]{4} }\)

OpenStudy (precalchelp):

got it thanks!

OpenStudy (solomonzelman):

can you say the answer you got, just in case...

OpenStudy (solomonzelman):

(No need, but if you want to...)

OpenStudy (precalchelp):

x^4√3 + 4√3x^4√3

OpenStudy (precalchelp):

woah oops hold on a sec

OpenStudy (solomonzelman):